# RB-SCD: A New Benchmark for Semantic Change Detection of Roads and Bridges in Traffic Scenes

### v2

@@ Abstract
With the rapid modernization of urban transportation, the accurate detection of changes such as construction, renovation, and demolition of roads and bridges plays a crucial role in urban planning and traffic management. However, existing methods often struggle to precisely extract semantic change information in such scenarios. A major limitation is the scarcity of fine-grained annotated change detection (CD) datasets in traffic scenes, which hinders the training and optimization of relevant algorithms. To address this, we introduce Road and Bridge Semantic Change Detection (RB-SCD) dataset, a comprehensive benchmark containing 260 pairs of high-resolution remote sensing images. The dataset spans a wide range of geographic regions and includes diverse road and bridge types across more than ten cities in multiple countries, incorporating 11 distinct categories of semantic changes, which collectively enable detailed structural and functional change analysis. Leveraging the rich diversity and complexity of RB-SCD, we further propose a novel framework, Multimodal Frequency-Driven Change Detector (MFDCD), which integrates multiple characteristics of multimodal features in the frequency domain for the first time. Technically, MFDCD consists of two components: Dynamic Frequency Coupler (DFC) and the Textual Frequency Filter (TFF). The DFC deeply couples hierarchical visual features with wavelet-derived frequency components, thereby enhancing the perception of detailed and cross-temporal global structural changes. The TFF converts textual features extracted by the pre-trained CLIP model into the frequency domain via Fourier transform. Then it applies graph-based filtering to extract salient frequency components, which are fused with visual features to enable joint multimodal representation learning. Experimental results show that MFDCD performs well on RB-SCD and three public datasets. The RB-SCD dataset, with its rich semantic annotations, serves as a challenging yet valuable resource for road and bridge change analysis in traffic environments. The code and dataset will be available at https://github.com/DaGuangDaGuang/RB-SCD.
Index Terms:
@@
Change detection (CD), road and bridge semantic change detection dataset, dynamic frequency coupler, textual frequency filter, remote sensing (RS), semantic map guidance.
I Introduction
@@
In the field of remote sensing (RS), change detection (CD) has become an essential task for monitoring environmental, urban, and infrastructure changes [1, 2]. With the acceleration of urbanization and the continuous development of transportation infrastructure, CD in traffic scenes is crucial for traffic planning and facility management. In traffic planning, accurately identifying structural and functional changes in roads and bridges helps urban authorities to assess the current state of transportation networks, forecast future demands, and make data-driven decisions about expansions, detours, and upgrades [3][4][5]. For instance, the emergence of new overpasses, flyovers, or road expansions needs to be timely integrated into transportation maps to ensure up-to-date routing and navigation services. Moreover, detecting illegal road constructions or temporary blockages is also critical for enforcing urban regulations and minimizing congestion. From the perspective of traffic facility management, road and bridge CD enables timely identification of structural deterioration, unauthorized demolitions, or unreported construction activities. This semantic-level understanding is crucial for preventive maintenance, disaster response (e.g., after floods or earthquakes), and resource allocation. Additionally, as smart cities and autonomous transportation systems increasingly rely on accurate and up-to-date digital maps, the need for automated and fine-grained CD becomes even more pressing. Roads and bridges are not only physical assets but also key elements in the digital infrastructure of intelligent transportation systems.
@@
However, as shown in Tab. I, traditional binary change detection (BCD) datasets [6][7][8], etc. primarily focus on building CD and only provide simple information about whether a change has occurred, represented by 0 and 1 for no change and change, respectively. On the other hand, existing semantic change detection (SCD) datasets [9][10][11], etc. mainly focus on changes in land cover, such as building changes and vegetation changes. Changes related to roads and bridges in traffic scenes are either limited in number or completely absent. Most existing road extraction datasets [12][13][14] mainly focus on binary pixel segmentation of roads and cannot be used to focus on road changes. To address this issue, as annotation examples shown in Fig. 1, we propose a new dataset specifically designed for SCD of roads and bridges in traffic scenes: Road and Bridge Semantic Change Detection (RB-SCD) dataset. RB-SCD provides semantic annotations for 11 distinct categories of road and bridge changes, effectively addressing insufficient granularity of information and lack of specificity in detecting road and bridge changes in traffic scenes.
In addition to the RB-SCD dataset, we have also developed a new framework to further facilitate CD performance, the Multimodal Frequency-Driven Change Detector (MFDCD), which integrates multimodal features in the frequency domain. Considering CD methods, traditional approaches often rely on convolutional neural network architectures [15][16][17], which are insufficient to fully exploit the rich spatiotemporal information in images. Some studies [18][19][20] have incorporated transformers into the network for global modeling. However, these methods still face challenges in accurately capturing and effectively distinguishing complex textures and structural changes. Research has also shown that relying solely on single-modal data has inherent limitations. Recently, some methods [21][22][23] have introduced multimodal features to overcome the shortcomings of single modalities, such as combining visual and textual information or integrating lidar and hyperspectral data. However, these approaches often fail to fully exploit the characteristics of multimodal features, focusing only on modeling the commonalities of multimodal data while neglecting the distributional differences among different modalities.
@@
TABLE I: Public remote sensing change detection and road extraction datasets. For image counting, the unit of change detection datasets is pairs, while that of road extraction is pieces/images.
@@
OSCD  [24] 10 24 $600\times 600$ Binary Land cover
S2MLTCP  [25] 10 1520 $600\times 600$ Binary Land cover
ZY3  [26] 5.8 1 $1154\times 740$ Binary Land cover
WV3  [26] 2 2 $1431\times 1431$ Binary Land cover
QB  [26] 2.4 1 $1154\times 740$ Binary Land cover
AICD  [27] 0.5 1000 $800\times 600$ Binary Land cover
@@
Google Earth  [14] 1.2 224 $512\times 512$ Binary Road
RB-SCD (Ours) 0.59 260 $270\times 534$ $7215\times 4366$ Semantic Roads and bridges 8
Figure 1: Annotation examples on RB-SCD dataset. $T_{1}$ and $T_{2}$ represent the pre- and post-change images, respectively. Label is the ground truth.
Figure 2: Visualization of visual and textual frequency spectrum.
In this work, as shown in Fig. 2, we observe that the decomposition and utilization of features in the frequency domain remain underexplored, particularly the fine-grained change information embedded in the high-frequency visual features, the large-scale distribution information in the low-frequency features, and the frequency-domain representation of textual features. The potential patterns of multimodal information in the frequency domain have not been adequately studied. Therefore, we make the first attempt to develop a new architecture that maps image and textual features into different types of frequency domains based on their characteristics, enabling multimodal frequency-domain modeling for CD.
Our MFDCD consists of two main components: Dynamic Frequency Coupler (DFC) and the Textual Frequency Filter (TFF). DFC decomposes bi-temporal features into high- and low-frequency components using discrete wavelet transform (DWT), and introduces an Adaptive Sparse Frequency Fusion (ASFF) block for the combination of high-frequency information and low-level features. During the early stages of training, ASFF enhances the sparsity of high-frequency features, focusing on significant change areas. In the later stages, the sparsity constraint is reduced to enhance the expressive power of the model, ultimately increasing the attention to key change areas while suppressing high-frequency noise interference. In this way, edge and texture information in low-level features can not only be gradually captured through convolution operations but also dynamically weighted based on the localized changes in refined high-frequency information. For low-frequency features, Bidirectional Temporal Frequency Fusion (BTFF) block is introduced. Since CD involves not only capturing local changes but also understanding the relative changes between different time points, simply combining the high-level features at a single time point with low-frequency features could cause the model to overly rely on the local structural information at that moment, neglecting the cross-temporal changes. The BTFF block addresses this by combining high-level features with low-frequency features across time points, enabling the model to understand the trend of changes and global structural changes over time.
TFF introduces a pre-trained CLIP [31] model to obtain textual information, which is further expanded to get text descriptors. Text features contain semantic information but lack spatial features, and discrete Fourier transform (DFT) can convert them into frequency domain information, making it easier to align with the frequency information of image features. By constructing graph filters to aggregate key information, the processed textual features are deeply integrated with visual features to provide multimodal support for SCD.
Our contributions are as follows:
  • A comprehensive benchmark for SCD in traffic scenarios. We construct a new dataset named RB-SCD, which focuses on SCD of roads and bridges. It contains 260 large-scale high-resolution RS image pairs, covering over ten cities across multiple countries and annotating 11 typical semantic change types. With its wide geographic diversity and complex traffic infrastructures, RB-SCD supports fine-grained structural and functional change analysis for transportation environments.
  • A novel multimodal frequency-driven CD framework. We propose Multimodal Frequency-Driven Change Detector (MFDCD), the first framework to introduce frequency-domain multimodal feature integration for CD. It consists of two main modules: (1) Dynamic Frequency Coupler (DFC), which fuses hierarchical visual features with high- and low-frequency components obtained via wavelet transform to enhance cross-temporal structural change perception; (2) Textual Frequency Filter (TFF), which transforms textual features extracted by a pre-trained CLIP model into the frequency domain using Fourier transform, applies graph-based filtering to extract salient components, and integrates them with visual features for collaborative representation learning.
@@
  • Comprehensive experiments on multiple benchmarks. Extensive experiments on the proposed RB-SCD and three public CD datasets demonstrate that MFDCD achieves superior performance and generalization ability in identifying complex semantic changes in traffic scenes.
II Related Work
II-A CD Datasets and Methods
@@
Existing datasets, as shown in Tab. I, such as  [6][7][8][28], are primarily designed for BCD tasks and cannot provide specific semantic change information. For SCD, datasets like  [9][10][11] have established some benchmarks. However, these datasets mainly focus on changes in land cover, with little attention paid to the changes in roads and bridges in traffic scenes. In contrast, our RB-SCD dataset is a high-resolution, fine-grained SCD dataset specifically curated and annotated for road and bridge changes in traffic scenes. It captures a wide range of semantic change types and includes diverse geographical contexts, enabling precise and meaningful evaluation of CD algorithms in real-world urban transportation scenarios. By addressing both the lack of semantic detail and the absence of domain specificity in existing datasets, RB-SCD provides a valuable and targeted benchmark for infrastructure monitoring, intelligent transportation analysis, and traffic scene understanding.
Figure 3: (a) Statistical distribution of pixels for semantic change categories. (b) Train set. (c) Test set.
TABLE II: Specific subcategories in RB-SCD dataset.
@@
Road street, highway, dirt road, walkway, railway, etc
@@
beam bridge, arch bridge, cable-stayed bridge,
suspension bridge, etc.
@@
wasteland, sand patches area, unpaved open
grounds, road or railway construction zones, etc.
@@
urban lawn, green belt, wetland grass, riparian
grass zone, meadow etc.
@@
Farmland agricultural land (with and without crops)
@@
residential building, commercial building,
industrial building, under-construction building
@@
In recent years, deep learning has significantly improved the accuracy of CD through its ability to extract relevant features from raw data. Changemask [32] decouples SCD into temporal semantic segmentation and BCD, and combines these two tasks into a unified encoder-transformer-decoder framework. MDENet [33] uses MDCM to capture object features in frequency and spatial domains across times, and MFGatt to reduce style impact, focusing on object location. JFRNet [34] introduces a three-branch network that takes joint bi-temporal features as unified input, enabling each branch to perceive the semantic relevance of the two time points without any additional interaction operations. DecoderFusion [35] enhances CD decoding features by leveraging the feature differences between two semantic segmentation decoding branches. CdSC [36] introduces a cross-differential semantic consistency network that explores the differences in bi-temporal features while maintaining their semantic consistency. HGINet [37] builds a multi-level perception aggregation network with a pyramid architecture to extract features that distinguish different categories across multiple levels. However, these methods only consider single-modal data, i.e., images, while neglecting the rich semantic information embedded in multimodal data, thus encountering a performance bottleneck.
II-B Multimodal Learning
Recent studies have shown that multimodal information is beneficial for visual image analysis. LAVT [21] and VLT [38] integrate textual features into visual features through language encoders. Zhang et al. [22] propose a convolutional neural network method with an attention mechanism to enhance feature extraction for lidar data. Their approach leverages the elevation information from lidar data and the spatial-spectral information of hyperspectral data for multimodal feature fusion. MedCLIP [39] uses text prompts to generate segmentation for clinical scans, adopting a novel decoupled hard-negative noise contrastive estimation loss to fine-tune the BiomedCLIP model. ChangeCLIP [23] designs a novel CD framework that utilizes semantic information from image-text pairs and introduce an innovative differential feature compensation module to capture fine-grained semantic changes between them.
II-C Frequency Transform in Deep Learning
In image processing, the combination of CNN and frequency domain information has been widely applied in tasks such as image restoration, super-resolution, and semantic segmentation. W-Unet [40] replaces down-sampling operations with DWT to reduce the impact of image noise while avoiding information loss. FFT-ReLU  [41] performs image deblurring tasks in the frequency domain based on fast Fourier transform. ARFFT [42] proposes a spatial-frequency fusion block to enhance the representation capability of Transformers and extend the receptive field across the entire image. XNet [43] utilizes wavelet transforms to transfer visual information to the frequency domain, leveraging high- and low-frequency information to design a framework supporting both fully supervised and semi-supervised tasks. Given the characteristics of RS images, such as high resolution and susceptibility to spectral variations caused by seasonal changes, using only frequency domain features often results in the loss of critical spatial information. Therefore, in this paper, we propose a novel architecture that models multimodal features in the frequency domain, effectively balancing the advantages of visual features, textual features, and frequency domain features, achieving promising results.
Figure 4: Examples of dataset distribution spanning multiple representative geographical areas.
@@ II-C Frequency Transform in Deep Learning
Figure 5: Examples of road and bridge types, where the first two columns are roads, and the last three columns are bridges.
III RB-SCD Dataset
III-A Data Annotation and Analysis
Rich Semantic Annotations: As shown in Fig. 1 and Fig. 3, unlike traditional BCD datasets, we have defined eight major categories (with specific subcategories detailed in Tab. II), including road, bridge, bare land, building, farmland, water, grass and vegetation. The RB-SCD dataset provides detailed semantic change annotations for 11 distinct types of semantic changes and a background category across these eight categories.
Figure 6: Overview of our proposed Multimodal Frequency-Driven Change Detector (MFDCD). Firstly, the Dynamic Frequency Coupler (DFC) couples hierarchical visual features extracted by ResNet50 with high- and low-frequency components obtained via wavelet transform to enhance cross-temporal structural change perception. Meanwhile, the Textual Frequency Filter (TFF) transforms textual features extracted by a pre-trained CLIP model into the frequency domain using Fourier transform, applies graph-based filtering to extract salient components, and integrates them with visual features for collaborative representation learning. Finally, encoder features are fed into the decoder to obtain the final semantic change detection results.
@@ III-A Data Annotation and Analysis
Changes such as road widening, new road construction, and bridge construction or modification are reflected through the corresponding semantic change types, supporting in-depth exploration of the intrinsic mechanisms of traffic scenario evolution. It is worth noting that some changes, such as bridges directly turning into roads, are rarely observed in reality. Additionally, other types of changes, such as those between bridges and some other categories, occur in only a few samples, and we set their pixel number to 0. All annotations are carried out by a professional team in RS and transportation, and underwent multiple rounds of rigorous review and quality control to ensure annotation accuracy and semantic consistency.
@@
Diverse Geographic Coverage: As shown in Fig. 4, following the principles in [44], we collect 260 pairs of RS images with resolutions ranging from 270 × 534 to 7,215 × 4,366 (0.59 m/pixel), which can be cropped into 5,016 pairs of 256 × 256 patches, including 4,015 pairs for the training set and 1,001 pairs for the testing set. The dataset integrates satellite images obtained primarily from Google Earth, covering various regions worldwide, including Chinese cities such as Hefei, Xi’an, Chongqing and Guangzhou, as well as representative areas from Japan, Singapore, Nigeria, Italy, America, and Indonesia, etc. This extensive geographic coverage ensures that the dataset encompasses traffic scenes under varying climatic conditions, geological structures, and urban planning contexts, thereby guaranteeing its diversity and representativeness.
Multiple Road and Bridge Types: As shown in Fig. 5, the dataset encompasses various types of roads and bridges from different geographical regions and urban environments. The road categories include concrete and asphalt roads used for urban main roads, secondary roads, residential streets, and highways, comprehensively reflecting the common transportation routes found in cities. Additionally, the dataset contains elevated roads designed to alleviate congestion and improve traffic efficiency, which are typically found in densely populated or high-traffic urban areas. In terms of bridge structures, the dataset includes large cable-stayed bridges that span wide rivers or deep valleys, as well as girder bridges and pedestrian bridges used for short- to medium-span connections and pedestrian access. The dataset also includes other types of specialized roads and bridges, demonstrating a variety of engineering styles and applications. This richness in road and bridge types enables the dataset to represent complex traffic environments and provides robust support for training CD models across multiple scales and diverse scenarios.
III-B Label Imbalance Analysis and Challenges

### v3

# Semantic Change Detection of Roads and Bridges: A Fine-grained Dataset and Multimodal Frequency-driven Detector

@@ Abstract
Accurate detection of road and bridge changes is crucial for urban planning and transportation management, yet presents unique challenges for general change detection (CD). Key difficulties arise from maintaining the continuity of roads and bridges as linear structures and disambiguating visually similar land covers (e.g., road construction vs. bare land). Existing spatial-domain models struggle with these issues, further hindered by the lack of specialized, semantically rich datasets. To fill these gaps, we introduce the Road and Bridge Semantic Change Detection (RB-SCD) dataset. As the first benchmark to systematically target semantic change detection of roads and bridges, RB-SCD offers comprehensive fine-grained annotations for 11 semantic change categories. This enables a detailed analysis of traffic infrastructure evolution. Building on this, we propose a novel framework, the Multimodal Frequency-Driven Change Detector (MFDCD). MFDCD integrates multimodal features in the frequency domain through two key components: (1) the Dynamic Frequency Coupler (DFC), which leverages wavelet transform to decompose visual features, enabling it to robustly model the continuity of linear transitions; and (2) the Textual Frequency Filter (TFF), which encodes semantic priors into frequency-domain graphs and applies filter banks to align them with visual features, resolving semantic ambiguities. Experiments demonstrate the state-of-the-art performance of MFDCD on RB-SCD and three public CD datasets. The code will be available at https://github.com/DaGuangDaGuang/RB-SCD.
keywords:
@@
Road and bridge change detection, Dynamic frequency coupler, Textual frequency filter, Remote sensing
1 INTRODUCTION
@@
In the field remote sensing (RS), change detection (CD) has become an essential task for monitoring environmental, urban, and infrastructure changes [1] . With the acceleration of urbanization and the continuous development of transportation infrastructure, accurately identifying structural and functional changes in roads and bridges helps urban authorities to assess the current state of transportation networks, forecast future demands, and make data-driven decisions about expansions, detours, and upgrades. Moreover, road and bridge CD enables timely identification of structural deterioration, unauthorized demolitions, or unreported construction activities [2].
@@
However, as shown in Tab. 1, traditional binary change detection (BCD) datasets such as LEVIR [3], WHU [4], CDD [5], SYSU [6], merely indicate whether a change has occurred through a simple binary mask (0 for no change, 1 for change). They lack fine-grained semantic details. While existing semantic change detection (SCD) datasets like Landsat [7], HRSCD [8], SECOND [9] mainly focus on changes in land cover. Changes related to roads and bridges in traffic scenes are either limited in number or completely absent. Although existing road extraction datasets [10][11] mainly focus on binary pixel segmentation of roads, they cannot be used to focus on road changes.
Methodologically, detecting semantic changes in roads and bridges presents unique challenges that are not fully addressed by general-purpose CD methods. Two key difficulties stand out. First, unlike the discrete, monolithic changes of buildings common in datasets (first row of Fig. 1), road networks evolve as winding linear structures and their changes often involve complex transitions from multiple, heterogeneous land-cover types such as water, vegetation, and bare land simultaneously (second row of Fig. 1).
@@
Table 1: Comparison of public remote sensing change detection and road extraction datasets. For image counting, the unit for change detection datasets is pairs of images, whereas for road extraction datasets, it is individual image. Notably, no existing dataset systematically addresses road and bridge semantic changes.
@@
RB-SCD (Ours) 0.59 260 $270\times 534$ - $7215\times 4366$ Semantic Roads and bridges 8
Figure 1: Visual comparison of change characteristics across prominent CD datasets. (a) LEVIR-CD and (b) SYSU-CD primarily feature discrete, monolithic building changes, (c) SECOND often focus on land-cover transitions. (d) RB-SCD introduces the unique challenges of detecting linear and heterogeneous change (source of change is heterogeneous) transitions, where winding roads are formed from multiple land-cover types. Furthermore, it presents high semantic ambiguity, with visually similar yet semantically distinct transitions such as water->road and vegetation->road, which are difficult for conventional models to distinguish.
The limited receptive fields of recent methods like SUNNet [13], ICIF [14], USSFCNet [15] prevent them from perceiving the overall structure of road networks. Consequently, while they may correctly identify numerous localized patches of change, they ultimately produce fragmented and disconnected change maps, failing to assemble these patches into a single, continuous linear feature. Standard attention models such as BIT [16], Changeformer [17], DMINet [18], ELGCNet [19] leverage global attention to overcome this locality, their standard mechanisms treat all spatial relationships equally, lacking the explicit inductive bias required to model heterogeneous, linear structures. Second, the task is plagued by high semantic ambiguity. Visual features alone are often insufficient to resolve ambiguities, such as distinguishing a road under construction from bare land (the first and third sample of Fig. 1 (d)), or differentiating shallow water from adjacent vegetation (second sample of Fig. 1 (d)). Even multimodal models like ChangeCLIP [20], which leverage vision-language for CD, may struggle when visual features themselves are highly ambiguous, as their fusion process operates within the noisy spatial domain.
Figure 2: Frequency-domain transformation of visual and textual data. (a) The DWT separates a visual scene into high-frequency (details) and low-frequency (structure) components. (b) The DFT converts a textual description into a global frequency spectrum, serving as a robust semantic signature for fusion.
The limitations of existing spatial-domain methods highlight the need for a new approach that operates in a more suitable domain. To robustly model linear structures, the core insight is that their essential continuity of structure is naturally encoded in low-frequency components, while sharp transition boundaries correspond to high-frequency components. This intrinsic property makes the frequency domain a logical choice. As shown in Fig. 2(a), we leverage the discrete wavelet transform (DWT) to decompose visual features and explicitly isolate these essential structural forms from fine-grained boundary details. To resolve deep semantic ambiguity, multimodal integration is required. However, prior methods suffer from fusing textual descriptions directly with unstable pixel-level features in the spatial domain. We instead transform text into its frequency spectrum via discrete Fourier transform (DFT) (Fig. 2(b)). This frequency-domain representation naturally filters out minor textual variations, providing a stable basis for a cleaner alignment with visual features and overcoming the core limitations of spatial-level fusion.
Based on these insights, we first introduce the Road and Bridge Semantic Change Detection (RB-SCD) dataset, a new benchmark specifically designed for SCD of roads and bridges. RB-SCD provides fine-grained semantic annotations for 11 distinct change categories across diverse geographic regions, enabling a detail evaluation of model capabilities in traffic domain. Second, we propose a novel Multimodal Frequency-Driven Change Detector (MFDCD). Our framework tackles the core challenges by operating in the frequency domain for both visual and textual features. It consists of two specialized components: (1) the Dynamic Frequency Coupler (DFC), which utilizes DWT operation to separate low-frequency components (encoding structural continuity) from high-frequency details (capturing transition boundaries). Within the DFC, an Adaptive Sparse Frequency Fusion (ASFF) block delineates boundaries by dynamically combining high-frequency information with low-level visual features. A Bidirectional Temporal Frequency Fusion (BTFF) block fuses low-level frequency and high-level visual features across time points to model the linear structure of change regions, preventing fragmentation; and (2) the Textual Frequency Filter (TFF), which first converts textual features into the frequency domain, a graph-based filtering mechanism then acts as a semantic reasoning to extract salient semantic components and joint multimodal representation learning with visual features. The resulting unambiguous features then guide the model toward confident semantic predictions. Our contributions are as follows:
  • 1.
    We introduce RB-SCD, a challenging benchmark dataset specifically designed for SCD of roads and bridges. It addresses the critical scarcity of domain-specific data by providing 11 fine-grained semantic change categories, enabling a more nuanced and practical analysis of traffic infrastructure evolution.
  • 2.
    We propose MFDCD, a framework that for the first time integrates multimodal (visual-textual) features in the frequency domain for CD. The framework overcomes conventional limitations through two key components: the DFC leverages wavelet transform to robustly model linear and heterogeneous semantic changes, and the TFF acts as a semantic reasoning engine to resolve high semantic ambiguity by aligning textual concepts with visual patterns in the frequency domain.
  • 3.
@@
    Extensive experiments show the state-of-the-art performance of our MFDCD. Specifically, on our proposed RB-SCD dataset, it excels at identifying complex semantic changes in traffic scenes. On three public benchmarks for BCD, it shows superior generalization ability and achieves highly competitive results.
2 Related Work
2.1 CD Datasets and Methods
@@
Existing BCD datasets such as LEVIR [3], WHU [4], CDD [5], SYSU [6] are primarily designed for BCD tasks and cannot provide specific semantic change information. For SCD, datasets like Landsat [7], HRSCD [8], SECOND [9] mainly focus on changes in land cover, with little attention paid to the changes in roads and bridges in traffic scenes. In contrast, our RB-SCD dataset is a high-resolution, fine-grained SCD dataset specifically curated and annotated for road and bridge changes in traffic scenes. It enables precise and meaningful evaluation of CD algorithms in traffic scenarios.
In recent years, deep learning has significantly improved the accuracy of CD through its ability to extract relevant features from raw data. C-3PO [21] reframes CD as a semantic segmentation task, proposing a new paradigm that adapts existing powerful segmentation networks for CD by introducing a dedicated module to learn different change types. S²PNet [22] enhances CD in complex VHR scenes by introducing a self-structured feature pyramid for better multi-scale object cognition. MHF²-Net [23] addresses the multi-scale object detection challenge in VHR CD by proposing a hierarchical feature fusion network that explicitly enhances high-frequency details. AGFormer [24] introduces a novel transformer-based model that mitigates classifier bias from class imbalance in CD by using an anchor-guided regularization strategy on a hypersphere to ensure better inter-class separability and intra-class balance. CdSC [25] introduces a cross-differential semantic consistency network that explores the differences in bi-temporal features while maintaining their semantic consistency. HGINet [26] builds a multi-level perception aggregation network with a pyramid architecture to extract features that distinguish different categories across multiple levels. However, these methods only consider single-modal data, i.e., images, while neglecting the rich semantic information embedded in multimodal data, thus encountering a performance bottleneck.
2.2 Multimodal Learning for Change Detection
Recent studies have shown that multimodal information is beneficial for visual image analysis. ChangeCLIP [20] designs a novel CD framework that utilizes semantic information from image-text pairs and introduce an innovative differential feature compensation module to capture fine-grained semantic changes between them. HFA-PANet [27] addresses the domain shift challenge in multimodal CD by proposing a non-siamese network that hierarchically aligns features from heterogeneous images to effectively capture their differences. AEKAN [28] introduces a novel unsupervised approach for multimodal CD by using a Siamese Kolmogorov-Arnold Network (KAN) AutoEncoder to effectively learn the latent commonality features between heterogeneous images. CFRL [29] presents an unsupervised framework for multimodal CD that learns comparable commonality features by cyclically reconstructing, re-encoding, and aligning features from heterogeneous images. While these pioneering works highlight the growing importance of multimodality, they primarily fall into two categories: visual-textual fusion in the spatial domain or feature alignment for heterogeneous image inputs. Our work carves out a distinct niche by being the first to introduce a visual-textual paradigm that operates within the frequency domain, specifically designed to resolve the deep semantic ambiguities found in complex, heterogeneous RS scenes
2.3 Frequency-Domain Analysis in Deep Learning
In image processing, the combination of CNN and frequency domain information has been widely applied in tasks such as image restoration, super-resolution, and semantic segmentation. W-Unet [30] replaces down-sampling operations with DWT to reduce the impact of image noise while avoiding information loss. FFT-ReLU  [31] performs image deblurring tasks in the frequency domain based on fast Fourier transform. ARFFT [32] proposes a spatial-frequency fusion block to enhance the representation capability of Transformers and extend the receptive field across the entire image. XNet [33] utilizes high- and low-frequency information to design a framework supporting both fully supervised and semi-supervised tasks.
Figure 3: Annotation examples on RB-SCD dataset. $T_{1}$ and $T_{2}$ represent the pre- and post-change images, respectively. The Label row shows the ground truth.
Given the characteristics of RS images, such as high resolution and susceptibility to spectral variations caused by seasonal changes, using only frequency domain features often results in the loss of critical spatial information. Therefore, in this paper, we model multimodal features in the frequency domain, effectively balancing the advantages of visual features, textual features, and frequency domain features, achieving promising results.
3 RB-SCD Dataset
As established, existing benchmarks are not suited for the task of road and bridge CD in traffic scenes. Therefore, our primary goal in creating the RB-SCD dataset is to provide detailed, fine-grained annotations that explicitly capture the semantic evolution of linear networks, while also covering a wide range of complex real-world scenes.
3.1 Data Annotation and Analysis
Rich Semantic Annotations. As shown in Fig. 3 and Fig. 4, unlike traditional BCD datasets, RB-SCD provides fine-grained annotations for 11 distinct semantic change types. These change annotations are constructed from eight fundamental land-cover classes, which are detailed in Tab. 2. A critical contribution of RB-SCD is its fine-grained semantic granularity, which is not merely a detail but an essential feature for meaningful real-world applications. This granularity allows for a precise understanding of urban evolution. For instance, categories like farmland->road and vegetation->road directly track the impact of urbanization on
Figure 4: Statistical analysis of the RB-SCD dataset. (a) Per-class pixel distribution for semantic change categories. (b-c) Semantic transition matrices for the training and test sets, respectively, showing the pixel counts for each from -> to change type. Here, bl, wat, veg, bd, fl, bri, and gra denote abbreviations for bare land, water, vegetation, building, farmland, bridge, and grass, respectively.
Table 2: Specific subcategories of land-cover classes in RB-SCD dataset.
@@ 3.1 Data Annotation and Analysis
Road street, highway, dirt road, walkway, railway, etc.
@@
beam bridge, arch bridge, cable-stayed bridge, suspension bridge, etc.
@@
wasteland, sand patches area, unpaved open grounds, road or railway construction zones, etc.
@@
urban lawn, green belt, wetland grass, riparian grass zone, meadow etc.
@@
Farmland agricultural land (with and without crops).
@@
residential building, commercial building, industrial building, under-construction building
@@
Figure 5: The diverse geographic distribution of the RB-SCD dataset. The samples span multiple countries and continents, including varied urban and rural landscapes from China, Japan, America, Italy, Singapore, Australia, and Indonesia. This diversity ensures the dataset’s robustness and challenges models to generalize across different environmental conditions and infrastructure styles.
@@
Figure 6: Examples of road and bridge types, where the first four columns are roads, and the last three columns are bridges.
@@
ecological and agricultural land, providing critical data for environmental assessment. Conversely, changes such as road->grass or the rare but critical bridge->water capture infrastructure degradation or demolition, which is vital for safety monitoring and disaster response. Distinguishing between road widening (an implicit category captured through pixel-level changes) and building->road is crucial for urban renewal analysis and traffic management. All annotations are performed by a professional team in RS and underwent rigorous quality control to ensure accuracy and semantic consistency. It is worth noting that some changes, such as bridges directly turning into roads, are rarely observed in reality. Additionally, other types of changes, such as those between bridges and some other categories, occur in only a few samples, and we set their pixel number to 0 in the training process.
@@
Diverse Geographic Coverage. The image pairs are sourced from Google Earth, with a temporal coverage spanning from 2007 to 2022. As shown in Fig. 5, following the principles in [34], we collect 260 pairs of RS images with resolutions ranging from 270 × 534 to 7,215 × 4,366 (0.59 m/pixel), which can be cropped into 5,016 pairs of 256 × 256 patches, including 4,015 pairs for the training set and 1,001 pairs for the testing set. The dataset covers various regions worldwide, including Chinese cities such as Hefei, Xi’an, Chongqing and Guangzhou, as well as representative areas from Japan, Singapore, Nigeria, Italy, America, and Indonesia, etc. This extensive geographic coverage ensures that the dataset encompasses traffic scenes under varying climatic conditions, geological structures, and urban planning contexts, thereby guaranteeing its diversity and representativeness.
Multiple Road and Bridge Types. As shown in Fig. 6, the dataset covers diverse road and bridge types across varied regions and urban contexts. Road categories range from urban main roads, secondary streets, and highways to elevated roads designed for congestion relief, while bridge types include large cable-stayed bridges, girder bridges, and pedestrian bridges. This diversity captures different engineering styles and traffic environments, providing robust support for training CD models across multiple scales and diverse scenarios.
3.2 Dataset Limitations